# Log-concavity And The Multiplicative Properties of Restricted Partition Functions

### v1

@@ 1.2. Other Log-Concave Sequences
There are many examples of log-concave sequences that, in light of Theorem 1.1, are now known to be abundant. Log-concavity has been established for Stirling numbers of the first and second kind [34], the Hyperfibonacci and Hyperlucas numbers [39], the HyperPell and HyperPell-Lucas numbers [1], and the figurate and central figurate numbers [20, 19] to name a few. A more general class of sequences is proven log-concave by Sullivan [33] who considered sequences $\{\alpha(n)\}$ of the form $\alpha(n)=n^{a}\exp(cn^{b})$ for various values of $a,b$ and $c$.
@@
Proofs that particular restricted partition functions are log-concave or abundant have typically relied on knowing their asymptotic. The $k$-regular partition function was proven abundant in 2016 by Beckwith & Bessenrodt [4], and log-concavity was proven in 2019 by Craig & Pun [9]. The $k$-colored partition function was proven both log-concave and abundant in 2018 by Chern, Fu, & Tang [8]. Dawsey & Masri [27] proved the Andrews smallest parts function is log-concave and abundant in 2017. In 2022, the plane partition function was proven abundant by Heim, Neuhauser, & Tröger [21] who also conjectured log-concavity; this was later proven that same year by Ono, Pujahari, & Rolen [31].
There are restricted partition functions for which log-concavity has been established, but which have not been proven to be abundant. Proofs of log-concavity for such functions are given by Dong & Ji [11] for the distinct partition function (equivalent to the partitions into odd parts function), by Engel [13] for the overpartition function, by Iskander, Jain, & Talvola [23] for the fractional partition function, by Gajdzica [15] for the multiset partition function, and independently by Benfield, Roy, & Paul [5] and by Sullivan [32] for the $k$-th power partition function.
This paper proves that these restricted partition functions are abundant. Where possible for these restricted partition functions, in the spirit of the original publication by Bessenrodt & Ono [30], this paper also calculates the finitely many values where the strict abundant inequality fails or is equal.
2. Proof of Theorem 1.1 and 1.2
Let $\{x_{n}\}_{n\in\mathbb{N}}$ be a sequence that is log-concave for all $n>N$. For a fixed $n>N$, assume $x_{k}^{2}\geq x_{k-1}x_{k+1}$ where $k=N+1,N+2,\dots,n$. By log-concavity, consider first the two inequalities $x_{n}^{2}\geq x_{n-1}x_{n+1}$ and $x_{n-1}^{2}\geq x_{n}x_{n-2}$. After some rearrangement, it follows that
@@
Again by log-concavity, $x_{n-2}^{2}\geq x_{n-1}x_{n-3}$, which can be substituted into (4), yielding
@@
Iterating this process yields the following inequality:
@@
This process may continues recursively, and as $n$ gets large,
@@
To prove the right hand side of Theorem 1.1, it is necessary to establish relations for all entries in the sequence, in particular, for $x_{0},x_{1},\dots,x_{N}$. To achieve this, define the sequence $\{a_{n}\}$ recursively by $a_{n}=x_{n}$ for all $n\geq N+1$ and $1\leq a_{n-1}\leq\frac{(a_{n})^{2}}{a_{n+1}}$ for $n=1,2,\ldots,N+1$. In particular, one may chose $a_{n}=x_{n}$ for all $n\geq N+1$ and for $n\leq N$,
$a_{N}:=\frac{(a_{N+1})^{2}}{a_{N+2}}=\frac{(x_{N+1})^{2}}{x_{N+2}}$
$a_{N-1}:=\frac{(a_{N})^{2}}{a_{N+1}}=\frac{\left(\frac{(x_{N+1})^{2}}{x_{N+2}}\right)^{2}}{x_{N+1}}=\frac{(x_{N+1})^{3}}{(x_{N+2})^{2}}$
@@
$a_{N-N}=a_{0}:=\frac{(x_{N+1})^{N+2}}{(x_{N+2})^{N+1}}.$
Hence, $(a_{n+1})^{2}\geq a_{n}a_{n+2}$ for all $n\geq 0$. A similar process to that outlined in (5) yields
@@
Note by Condition (2) that $a_{0}\geq 1$. And by log-concavity, $a_{1}^{2}\geq a_{0}a_{2}\geq a_{2}$ implying that $\frac{a_{2}}{a_{1}}\leq a_{1}$. Then from (10) it follows that
@@
Hence
Both (13) and (14) cannot be simultaneously true, this gives a contradiction. Hence $a_{m+n}\leq a_{m}a_{n}$ for all $m,n\geq 0$ and in particular $m,n\geq N+1$ completing the proof of Theorem 1.1. ∎
Consider the consequences of starting this proof with the sequence $y_{n}=x_{n}^{-1}$. Then $y_{n}$ is log-convex for all $n>N$ and the same proof can be adapted where the only change is the reversal of every inequality sign, completing the proof of Theorem 1.2. ∎
@@ 3. Various Partition Functions
This section considers restricted partition functions $P_{A}(n)$ that are known to be log-concave. It follows from Theorem 1.1 that these sequences are also abundant, provided they also satisfy Condition (2). This section considers a few, well-studied sequences and partition functions and determines for which small indices abundance fails.
3.1. The Distinct Partition Function, Partitions into Odd Parts
The Distinct Partition Function $P_{d}(n)$ counts the number of ways to partition $n$ into distinct parts, where for all $i$ and $j$, $\lambda_{i}\neq\lambda_{j}$.
[…]
Curiously, Euler showed that $P_{d}(n)$ is equivalent to the number of partitions of $n$ using only odd integers [14, 25, 28]. For example, $P_{d}(4)=2$ since there are only two ways to partition $4$ into distinct parts: $4=4=3+1$ (and only two ways to partition $4$ into odd parts: $4=3+1=1+1+1+1$). In 2023, Dong & Ji [11] proved that the sequence $\{P_{d}(n)\}$ is log-concave for all $n\geq 33$. Note that $P_{d}(n)$ satisfies Condition (2) for $n=33$: $448^{\frac{1}{33}}>512^{\frac{1}{34}}$. The result follows from Theorem 1.1 and numerical computation for small indices.
Theorem 3.1.
The distinct partition function satisfies the inequality $P_{d}(a)P_{d}(b)\geq P_{d}(a+b)$ where instances of equality and failure occur as given in Table 1.
Table 1. Equality and Failure for $P_{d}(a)P_{d}(b)\geq P_{d}(a+b)$
Equality at $(a,b)$ Failure at $(a,b)$
$(1,1),(3,1),$ $a=1,2$
$a=5,6,7,8\quad b=3,$ $(4,3),(5,5)$
$a=15,16,17\quad b=4,$ $44\leq a\leq 14\quad b=$ 4
$a=6,7,8\quad b=5$
3.2. The Overpartition Function
The overpartition function, $\overline{p}(n)$, counts the number of partitions of $n$ in which the last occurrence of a number can be distinguished. For example, $\overline{p}(4)=14$:
[…]
In 2014, Engel [13] proved that $\{\overline{p}(n)\}$ function is log-concave for $n\geq 2$. Note that $\overline{p}$ satisfies Condition (2) for $n=2$: $\sqrt{4}\geq\sqrt[3]{8}$. The result follows from Theorem 1.1 and numerical computation for small indices.
Theorem 3.2.
The overpartition function satisfies the inequality $\overline{p}(a)\overline{p}(b)\geq\overline{p}(a+b)$ where instances of equality occur at $(1,1)$ and $(1,2)$.
The following subsections, 3.3, 3.4, and 3.5, have been proven to satisfy log-concavity for sufficiently large $n$, however, the smallest such index $N$ such that each sequence is log-concave for all $n\geq N$ is either unknown or only conjectured. On the other hand, it is known that the asymptotics of these sequences are each of the form $n^{a}e^{c\cdot n^{b}}$ for real numbers $a,b,$ and $c$ where $0<b\leq 1$. This implies that the $n$th root of these sequences are each decreasing, hence, there is an index $N$ such that $x_{N+1}^{\frac{1}{N+1}}\geq x_{N+2}^{\frac{1}{N+2}}$.
3.3. Fractional Partition Function
$\sum_{n=0}^{\infty}p_{\alpha}(n)q^{n}=(q;q)_{\infty}^{\alpha}$
In 2019 Huat & Chan [6] studied the properties of $p_{\alpha}(n)$ where alpha is a rational number. The next year, Iskander, Jain & Talvola [23] proved that for any real $\alpha$ there exists an integer $N_{\alpha}$ such that the fractional partition function $p_{\alpha}(n)$ is log-concave $n>N_{\alpha}$.
Theorem 3.3.
There is an index $M_{\alpha}$ such that the sequence $p_{\alpha}(n)$ satisfies the inequality $p_{\alpha}(n)p_{\alpha}(m)>p_{\alpha}(n+m)$ for all $n,m>M_{\alpha}$
3.4. The Multiset Partition Function
Let $k$ be a fixed positive integer and let $\mathcal{A}=\{a_{i}\}_{i=1}^{\infty}$ be a non-decreasing sequence of positive integers and define a restricted partition of $n$ the sequence of positive integers $\lambda_{1},\lambda_{2},\ldots,\lambda_{j}$ such that $n=\lambda_{1}+\lambda_{2}+\ldots+\lambda_{j}$ where each $\lambda_{i}$ belongs to the multiset $\{a_{1},a_{2},...,a_{k}\}$. The multiset partition function $p_{\mathcal{A}}(n,k)$ counts all partitions of $n$ among the first $k$ elements of the multiset $\mathcal{A}$. The sequence $\{p_{\mathcal{A}}(n)\}$ was studied in 2022 by Gajdzica [15] who proved log-concavity under certain conditions of $n$ and $k$. This result was futher improved by Gajdzica, Heim, & Markus [16].
Theorem 3.4.
There exists a sufficiently large index such that the multiset partition function satisfies the inequality $p_{\mathcal{A}}(n)p_{\mathcal{A}}(m)\geq p_{\mathcal{A}}(n+m)$ under the same conditions Gajdzica gives in his Theorem 1.1 [15].
@@ 3.5. The $k$-th Power Partition Function
Independently, both Benfield, Paul, & Roy [5] and Sullivan [32] proved that $p^{k}(n)$ is log-concave for all $k$ and sufficiently large $n$.
@@
There exists an index $M_{k}$ such that for all $k$, the inequality $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$ holds for $a,b>M_{k}+1$.
@@
Table 2. Smallest $N_{k}$ for $k=1,\ldots,6$, OEIS: A346160 [37]
$p^{k}(n)$ $p^{1}(n)$ $p^{2}(n)$ $p^{3}(n)$ $p^{4}(n)$ $p^{5}(n)$ $p^{6}(n)$
$N_{k}$ $25$ $1041$ $15655$ $637854$ $2507860$ $35577568$
Conjecture 3.7 gives instances $(a,b)$ where the inequality $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$ is equal and when it fails for $k=1,2,3$. The case $k=1$ was proven by Bessenrodt & Ono [30]. Note that equality always holds for $a+b<2^{k}$.
Conjecture 3.7.
The $k$-th power partition function satisfies the inequality $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$ with instances of failure or equality as given in Table 3:
Table 3. Equality and Failure for $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$
$p_{k}(n)$ Equality Failure
$a<2^{1},\ a+b>2^{1}$
$p_{1}(n)$ $(2,7),(2,6),(3,4)$ $a=2,3,4,5\quad b=2$
$a=3,5\quad b=3$
$a=5,6,7\quad b=4$ $a<2^{2},\ a+b>2^{2}$
$p_{2}(n)$ $a=5,6,8,11,15\quad b=5$ $(7,5)$
$a=8,10,12,14\quad b=6$ $a=6,7,11,15\quad b=6$
$a=8,9,12,13,19\quad b=7$ $a=7,10,11,14,15\quad b=7$
$a=15,23\quad b=9$ $a<2^{3},\ a+b>2^{3}$
$p_{3}(n)$ $a=14,15,22,23\quad b=10$ $a=15,23\quad b=12$
$a=13,14,15,21,22,23\quad b=11$ $a=14,15,22,23,b=13$
$a=12,13,14,20,21,22\quad b=12$ $a=14,15,21,22,23\quad b=14$
$a=13,19,20,21,23\quad b=13$ $a=15,20,21,22,23\quad b=15$
$a=18,19,20,26\quad b=14$
$a=17,18,19,26\quad b=15$
$(23,20),(22,21),(23,21),(22,22),(23,22),(23,23)$
@@ 4. Log-polynomial Sequences
In a more general setting, Sullivan [33] made interesting connections between the data of log-polynomial sequences and log-concavity/log-convexity. A sequence $\alpha(n)$ is log-polynomial of degree $m\geq 2$ with data $\{A(n),\kappa,\delta(n)\}$ for $\kappa=\pm 1$ if there exists $\delta(n)>0$, $A(n)$, and $g_{k}(n)$ for $k=3,4,\ldots,m$ such that
$\log{\left(\frac{\alpha(n+j)}{\alpha(n)}\right)}=A(n)j+\kappa\cdot\delta(n)^{2}j^{2}+\sum_{k=3}^{m}g_{k}(n)j^{k}+o\left(\delta(n)^{m+1}\right)$
for $j=1,2,\ldots,m+1$ where both $\delta(n)\rightarrow 0$ and $g_{k}(n)=o\left(\delta(n)^{k}\right)$ as $n\rightarrow\infty$. Sullivan [33] showed that if $\alpha(n)$ is log-polynomial for degrees $m\geq 2$, then $\alpha(n)$ is log-concave for large $n$ whenever $\kappa=-1$ and is log-convex for large $n$ whenever $\kappa=1$. In particular, Sullivan showed [33] that if $\alpha(n)=n^{a}e^{c\cdot n^{b}}$ for real numbers $a$, $b$, and $c$ where $a$ and $c$ are not both zero, $20<b<$ 2, and $b(b-1)c\neq 0$, then $\alpha(n)$ is log-polynomial with data $\{A(n),\kappa,\delta(n)\}$ where
$A(n)=-\frac{a}{n}+bc\cdot n^{b-1}\quad\ \quad\kappa=\text{sgn}(b(b-1)c)\quad\ \quad\kappa\cdot\delta(n)^{2}=-\frac{a}{2n^{2}}+\frac{b(b-1)c}{2n^{2-b}}$
@@
If a sequence $\alpha(n)$ is log-polynomial for degrees $m\geq 2$ with data $\{A(n),\kappa,\delta(n)\}$ and $0<b\leq 1$, then for sufficiently large integers $n$ and $m$, the sequence satisfies $\alpha(n)\alpha(m)\geq\alpha(n+m)$ whenever $\kappa=-1$ and satisfies $\alpha(n)\alpha(m)\leq\alpha(n+m)$ whenever $\kappa=1$ with finitely many instances of equality or failure.

### v4

@@ 1.3. Other Log-Concave Sequences
There are many examples of log-concave sequences that, in light of Theorem 1.1, are now known to be abundant. Log-concavity has been established for Stirling numbers of the first and second kind [37], the Hyperfibonacci and Hyperlucas numbers [42], the HyperPell and HyperPell-Lucas numbers [1], and the figurate and central figurate numbers [21, 20] to name a few. A more general class of sequences is proven log-concave by O’Sullivan [36] who considered sequences $\{a(n)\}$ of the form $a(n)=n^{r}\exp(tn^{s})$ for various values of $r,s$ and $t$.
@@
Proofs that particular restricted partition functions are log-concave or abundant or both have typically relied on knowing their asymptotic. The $k$-regular partition function for $2\leq k\leq 6$ was proven abundant in 2016 by Beckwith & Bessenrodt [4], and log-concavity was proven in 2019 by Craig & Pun [11] for large $n$. Proofs of log-concavity for the $k$-regular partition with $k=2$ and $3$ are given by Dong & Ji [13]. The $k$-colored partition function was proven abundant in 2018 by Chern, Fu, & Tang [10]. Log-concavity of the $k$-colored partition was proved by Bringmann, Kane, Rolen, & Tripp [7] which was conjectured in [10]. Dawsey & Masri [30] proved the Andrews smallest parts function is log-concave and abundant in 2017. In 2022, the plane partition function was proven abundant by Heim, Neuhauser, & Tröger [23] who also conjectured log-concavity; this was later proven that same year by Ono, Pujahari, & Rolen [34].
There are restricted partition functions for which log-concavity has been established, but which have not been proven to be abundant. Proofs of log-concavity for such functions are given by Iskander, Jain, & Talvola [25] for the fractional partition function, by Gajdzica [16] for the multiset partition function, and independently by Benfield, Roy, & Paul [5] and by O’Sullivan [35] for the $k$-th power partition function.
This paper proves that these restricted partition functions are abundant. Where possible for these restricted partition functions, in the spirit of the original publication by Bessenrodt and Ono [6], this paper also calculates the finitely many values where the strict abundant inequality fails or is equal.
2. Proof of Theorem 1.1
Let us assume $x_{k}^{2}\geq x_{k-1}x_{k+1}$ where $k=N+1,N+2,\dots,n$. Now consider the two inequalities $x_{n}^{2}\geq x_{n-1}x_{n+1}$ and $x_{n-1}^{2}\geq x_{n}x_{n-2}$. After some rearrangement, it follows that
@@
Since $x_{n-2}^{2}\geq x_{n-1}x_{n-3}$, substituting this into (9), yielding
@@
Iterating this process yields the following chain of inequalities:
@@
This process may continue recursively, and as $n$ gets large,
@@
To prove the abundantness, it is necessary to establish relations for all entries in the sequence. To achieve this, construct a sequence $\{a_{n}\}$ recursively by $a_{n}=x_{n}$ for all $n\geq N+k$ and for $n<N+k$,
$a_{N+k-1}:=\frac{(a_{N+k})^{2}}{a_{N+k+1}}=\frac{(x_{N+k})^{2}}{x_{N+k+1}}$ (15)
$a_{N+k-2}:=\frac{(a_{N+k-1})^{2}}{a_{N+k}}=\frac{\left(\frac{(x_{N+k})^{2}}{x_{N+k+1}}\right)^{2}}{x_{N+k}}=\frac{(x_{N+k})^{3}}{(x_{N+k+1})^{2}}$
@@
$a_{0}=\frac{(x_{N+k})^{N+k+1}}{(x_{N+k+1})^{N+k}}.$ (16)
Hence, $(a_{n+1})^{2}\geq a_{n}a_{n+2}$ for all $n\geq 0$ and $a_{0}\geq 1$ by Condition (3). Now from Asai, Kubo, & Kuo [3, Theorem 2] or Gajdzica, Miska, & Ulas [18, Theorem 4.4], one obtains
$a_{n}a_{m}\geq a_{n+m}\quad\text{for all}\quad m,n\geq 0.$ (17)
Here, we present an alternative proof of (17) compared to the one given in [3, 18], as it aligns better with the approach used in the initial part of the proof. A similar process to that outlined in (10) yields
@@
Note by Conditions (3) and (16) that $a_{0}\geq 1$. And by log-concavity, $a_{1}^{2}\geq a_{0}a_{2}\geq a_{2}$ implying that $\frac{a_{2}}{a_{1}}\leq a_{1}$. Then from (18) it follows that
@@
Hence both (21) and (22) cannot be simultaneously true, this gives a contradiction. Hence $a_{m+n}\leq a_{m}a_{n}$ for all $m,n\geq 0$ and in particular $m,n\geq N+k$.
Now, since $M\geq N+k\geq N$ then following the same steps of the proof of (14) we find
$x_{n}\geq(x_{M})^{\frac{m}{n-M+m}}(x_{n+m})^{\frac{n-M}{n-M+m}}$
and in particular
$x_{M+1}\geq(x_{M})^{\frac{m}{m+1}}(x_{M+1+m})^{\frac{1}{m+1}}.$
This gives
$\frac{x_{M+1+m}}{x_{M+1}}\leq\left(\frac{x_{M+1}}{x_{M}}\right)^{m}.$
Therefore by Condition (4)
$\frac{x_{M+1+m}}{x_{M+1}}\leq x_{m}$
for all $d\leq m\leq N+k$. Since the sequence $\{\frac{x_{n+1}}{x_{n}}\}_{n\geq N}$ is decreasing then
$\frac{x_{n+m}}{x_{n}}\leq\frac{x_{M+1+m}}{x_{M+1}}\leq x_{m}$
for all $n\geq M+1$ and $d\leq m\leq N+k$. This completes the proof of Theorem 1.1.
@@ 3. Various Partition Functions
This section considers restricted partition functions $P_{A}(n)$ that are known to be log-concave. It follows from Theorem 1.1 that these sequences are also abundant, provided they also satisfy Condition (3). This section considers a few, well-studied sequences and partition functions and determines for which small indices abundance fails. The following subsections have been proven to satisfy log-concavity for sufficiently large $n$, however, the smallest such index $N$ such that each sequence is log-concave for all $n\geq N$ is either unknown or only conjectured. On the other hand, it is known that the asymptotics of these sequences are each of the form $n^{r}e^{t\cdot n^{s}}$ for real numbers $s,r,$ and $t$ where $0<s\leq 1$. This implies that for some $n\geq n_{0}$ the $n$th root of these sequences are each decreasing, hence, there is an index $N$ such that $x_{N+1}^{\frac{1}{N+1}}\geq x_{N+2}^{\frac{1}{N+2}}$.
@@ 3.1. The $k$-th Power Partition Function
Independently, both Benfield, Paul, & Roy [5] and O’Sullivan [35] proved that $p^{k}(n)$ is log-concave for all $k$ and sufficiently large $n$.
@@
There exists an index $N_{k}$ such that for all $k$, the inequality $p^{k}(m)p^{k}(n)\geq p^{k}(m+n)$ holds for $n\geq m\geq N_{k}$.
Conjecture 3.2 gives instances $(m,n)$ where the inequality $p^{2}(m)p^{2}(n)\geq p^{2}(m+n)$ is equal and when it fails. The case $k=1$ was proven by Bessenrodt & Ono [33].
Conjecture 3.2.
The square power partition function satisfies the inequality $p^{2}(m)p^{2}(n)\geq p^{2}(m+n)$ with instances of failure or equality as given in Table $2$:
Equality at $(m,n)$ with $n\geq m$ Failure at $(m,n)$ with $n\geq m$ $m=1$, $n=1,2,4,5,6,9,10,13,14,18,22$ $m=1$, $n\neq 1,2,4,5,6,9,10,13,14,18,22$ $m=2,n=4,5,9,13$ $m=2$, $n\neq 1,4,5,9,13$ $m=3$, $n=4$ $m=3$, $n\neq 1,2,4$ $m=4$, $n=5,6,7$ $m=5$, $n=5,6,8,11,15$ $m=5$, $n=7$ $m=6$, $n=8,10,12,14$ $m=6$, $n=6,7,11,15$ $m=7$, $n=8,9,12,13,19$ $m=7$, $n=7,10,11,14,15$
Table 2. Equality and Failure for $p^{2}(m)p^{2}(n)\geq p^{2}(m+n)$
@@
Table 3. Smallest $N_{k}$ for $k=2,\ldots,6$, OEIS: A346160 [40]
$p^{k}(n)$ $p^{2}(n)$ $p^{3}(n)$ $p^{4}(n)$ $p^{5}(n)$ $p^{6}(n)$
$N_{k}$ $1041$ $15655$ $637854$ $2507860$ $35577568$
3.2. Fractional Partition Function
$\sum_{n=0}^{\infty}p_{\alpha}(n)q^{n}=(q;q)_{\infty}^{\alpha}$
In 2019 Huat & Chan [8] studied the properties of $p_{\alpha}(n)$ where $\alpha$ is a rational number. The next year, Iskander, Jain, & Talvola [25] proved that for any real $\alpha$ there exists an integer $N_{\alpha}$ such that the fractional partition function $p_{\alpha}(n)$ is log-concave $n>N_{\alpha}$.
Theorem 3.4.
Let $\alpha>0$. Then there is an index $M_{\alpha}\geq N_{\alpha}$ such that the sequence $p_{\alpha}(n)$ satisfies the inequality $p_{\alpha}(n)p_{\alpha}(m)>p_{\alpha}(n+m)$ for all $n,m>M_{\alpha}$
3.3. The Multiset Partition Function
Let $k$ be a fixed positive integer and let $\mathcal{A}=\{a_{i}\}_{i=1}^{\infty}$ be a non-decreasing sequence of positive integers and define a restricted partition of $n$ as the sequence of positive integers $\lambda_{1},\lambda_{2},\ldots,\lambda_{j}$ such that $n=\lambda_{1}+\lambda_{2}+\ldots+\lambda_{j}$ where each $\lambda_{i}$ belongs to the multiset $\{a_{1},a_{2},...,a_{k}\}$. The multiset partition function $p_{\mathcal{A}}(n,k)$ counts all partitions of $n$ among the first $k$ elements of the multiset $\mathcal{A}$. The sequence $\{p_{\mathcal{A}}(n,k)\}$ was studied in 2022 by Gajdzica [16] who proved log-concavity under certain conditions of $n$ and $k$. This result was further improved by Gajdzica, Heim, & Neuhauser [17].
Theorem 3.5.
There exists a sufficiently large index such that the multiset partition function satisfies the inequality $p_{\mathcal{A}}(n,k)p_{\mathcal{A}}(m,k)\geq p_{\mathcal{A}}(n+m,k)$ under the same conditions Gajdzica gives in his Theorem 1.1 [16].
@@ 4. Log-polynomial Sequences
In a more general setting, O’Sullivan [36] made interesting connections between the data of log-polynomial sequences and log-concavity/log-convexity. A sequence $a(n)$ is log-polynomial of degree $m\geq 2$ with data $\{A(n),\kappa,\delta(n)\}$ for $\kappa=\pm 1$ if there exists $\delta(n)>0$, $A(n)$, and $g_{k}(n)$ for $k=3,4,\ldots,m$ such that
$\log{\left(\frac{a(n+j)}{a(n)}\right)}=A(n)j+\kappa\cdot\delta(n)^{2}j^{2}+\sum_{k=3}^{m}g_{k}(n)j^{k}+o\left(\delta(n)^{m+1}\right)$
for $j=1,2,\ldots,m+1$ where both $\delta(n)\rightarrow 0$ and $g_{k}(n)=o\left(\delta(n)^{k}\right)$ as $n\rightarrow\infty$. O’Sullivan [36] showed that if $a(n)$ is log-polynomial for degrees $m\geq 2$, then $a(n)$ is log-concave for large $n$ whenever $\kappa=-1$ and is log-convex for large $n$ whenever $\kappa=1$. In particular, he showed [36] that if $a(n)=n^{r}e^{t\cdot n^{s}}$ for real numbers $r$, $s$, and $t$ where $r$ and $t$ are not both zero, $20<s<$ 2, and $s(s-1)t\neq 0$, then $a(n)$ is log-polynomial with data $\{A(n),\kappa,\delta(n)\}$ where
$A(n)=-\frac{r}{n}+st\cdot n^{s-1}\quad\ \quad\kappa=\text{sgn}(s(s-1)t)\quad\ \quad\kappa\cdot\delta(n)^{2}=-\frac{r}{2n^{2}}+\frac{s(s-1)t}{2n^{2-s}}$
@@
If a sequence $a(n)$ is log-polynomial for degrees $m\geq 2$ with data $\{A(n),\kappa,\delta(n)\}$ and $0<s\leq 1$, then for sufficiently large integers $n$ and $m$, the sequence satisfies $a(n)a(m)\geq a(n+m)$ whenever $\kappa=-1$ and satisfies $a(n)a(m)\leq a(n+m)$ whenever $\kappa=1$ with finitely many instances of equality or failure.